CH8PT.EC

Morganton Company makes one product and it provided the following information to help prepare the master budget: a. The budge7. In July what are the total estimated cash disbursements for raw materials purchases? Assume the cost of raw material purch8. If 100,800 pounds of raw materials are needed to meet production in August, what is the estimated accounts payable balance9. If 100,800 pounds of raw materials are needed to meet production in August, what is the estimated raw materials inventory10. What is the total estimated direct labor cost for July assuming the direct labor workforce is adjusted to match the hours11. If we assume that there is no fixed manufacturing overhead and the variable manufacturing overhead is $8 per direct labor12. If we assume that there is no fixed manufacturing overhead and the variable manufacturing overhead is $8 per direct labor13. If we assume that there is no fixed manufacturing overhead and the variable manufacturing overhead is $8 per direct labor15. If we assume that there is no fixed manufacturing overhead and the variable manufacturing overhead is $8 per direct labor
Answer #1

7.total disbursement for the July = 70% of June and 30% of July=$(131040*70%+(92000*30%)=$119328

8.account payable at the end of July =70% July purchase(230000*.70) =$161000

9.closing stock raw materials =(next month requirement *10%)=10800*10%=1080 units

10.total labour cost for the july =(23000*2*13)=$598000

11. computation of average cost

raw materials (4*2.50) 10.00
direct labour (13*2) 26.00

variable manufacturing cost (8*2)

16.00
variable selling and administrative cost 1.80
Fixed selling and administrative cost (62000*4)/83200 2.98
average cost    56.78

12.value of finished goods at the end of July = 25000*.20*54.98=$274903.85

13. cost o goods sold for the month of July =25000*.80*(10+26+16)= $1040000

Gross margin =slaes -the cost of goods sold =(25000*60*.8)-1040000=$160000

14. total selling and administrative cost for the month of July= (23000*1.80 +62000)=$103400

15. net operating income = (25000*.80(60-56.78)=$64400